OpenStudy (anonymous):

sin 2x + sin x = 0 Find the exact solutions of the equation in the interval [0, 2π) Dont really understand this, anyone have any answers?

OpenStudy (anonymous):

\[\sin2x+\sin x=0\] or \[\sin^2x+\sin x=0\]

OpenStudy (anonymous):

is it \(\sin^2(x)\) or \(\sin(2x)\) ?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Gotta be sure :)

OpenStudy (anonymous):

This is a practice one and the answer sin 2x − sin x = 0 0, pie/3, pie, 5pie/3 I dont get how they got it

OpenStudy (anonymous):

That's not what I asked... Is the sine squared, or is the x multiplied by 2?

OpenStudy (anonymous):

mmmm pie \(\pi\)

OpenStudy (anonymous):

your top answer its sin 2x + sin x =0

OpenStudy (imstuck):

It can't be sin squared; the answers to that are pi, 2pi, and pi/2

mathslover (mathslover):

It is given to you that : \(\sin(2x) + \sin x = 0\) We can write that as : \(\sin(2x) = -\sin (x)\) Right?

OpenStudy (anonymous):

IMstuck I tried it and doesent work

OpenStudy (anonymous):

yeah that should work

mathslover (mathslover):

Zrander, as you tried to find the answer, I would love to see what you tried yet. And will be better if I recognize the mistake in your question.

OpenStudy (anonymous):

\[\sin2x+\sin x=0\\ 2\sin x\cos x+\sin x=0\\ \sin x(2\cos x+1)=0\] \[\sin x=0~~~~\text{and}~~~~\cos x=-\frac{1}{2}\]

OpenStudy (imstuck):

If you have sin(2x)= -sin(x), then what? where do you go with it? Even if you use the fact that sin(2x) = 2sinxcosx, you get c0s(x) = -1/2, not positive 1/2. I think something is wrong.

mathslover (mathslover):

Or you can just follow this : \(\sin (2x) = -\sin (x)\) \(2\sin (x) \cos (x) = - \sin (x) \) \(2\cos (x) = - 1\) or \(\sin (x) =0 \) \(\cos (x) = \cfrac{-1}{2} \) or \(\sin (x) = 0 \)

OpenStudy (anonymous):

Find the exact solutions of the equation in the interval [0, 2π). (Enter your answers as a comma-separated list.) sin 2x + sin x = 0 This is the question word for word. Idk. I just dont get it

OpenStudy (imstuck):

The cosine of x is -1/2 at 2pi/3 and 4pi/3, not pi/3 and 5pi/3 as zrander23 says they are. It just doesn't come out that way.

OpenStudy (imstuck):

the solutions to the problem above are 0, pi, 2pi/3, 4pi/3

OpenStudy (anonymous):

This is one of the practice ones on my homework, as i stated earlier, I was just showing how they wanted it formatted sin 2x − sin x = 0 0, pie/3, pie, 5pie/3

OpenStudy (imstuck):

Use the unit circle. Are you allowed to use a unit circle to find the answers once you solve for x?

OpenStudy (anonymous):

yeah

OpenStudy (imstuck):

Ok, I think I get it, what you're saying. You're saying that the 0, pi/3, pi, 5pi/3 are just EXAMPLES of the format of the answers you find?

OpenStudy (anonymous):

yes haha

OpenStudy (imstuck):

Ok. Then here we go with solving this mess! ; )

OpenStudy (anonymous):

Thoose asnwers you gave me worked, i somewhat understand now!

OpenStudy (anonymous):

I looked at my unit circle haha

OpenStudy (imstuck):

Do you want to understand thoroughly or do you just want to somewhat understand? I'm a high school math teacher and somewhat understanding does not usually work out too well from my experience! Let us help you if you need to understand better. Do you know how to solve the equations to find what x is?

OpenStudy (anonymous):

I give you credit for being a math teacher!, not really

OpenStudy (anonymous):

i dont know how to solve it

OpenStudy (anonymous):

I struggle with math and this is my last math class i need to talk for college, for my degree

OpenStudy (imstuck):

You want help? I would love to help you!

OpenStudy (anonymous):

Ill send you a message

OpenStudy (anonymous):

Yeah I mean I could use help, for my online class I have one more test and a final

OpenStudy (imstuck):

Ok the first thing you have to do is realize that sin(2x) is a double angle. In order to deal with it, you need to use its equivalency and get rid of it. Are you familiar with what the identity of sin(2x) is?

OpenStudy (anonymous):

2 (sin x )(cos x ) right?

OpenStudy (imstuck):

Yes. So just replace the sin(2x) with 2sinxcosx to get 2sin(x)cos(x) + sin(x) = 0. Now you have a situation where you can simplify by factoring. You see there is a sin(x) on both sides of the " + " sign, so you can factor one out. What do you get when you do that?

OpenStudy (imstuck):

2sin(x)cos(x) + sin(x) = 0 becomes \[\sin x(2\cos x +1)=0\]

OpenStudy (anonymous):

i was working it on paper and thats what i just got

OpenStudy (imstuck):

Ok, that is a perfect start! the sin(x) = 0 is an easy one. So let's deal with the 2 cos(x) + 1 = 0 first. You solve this like any "normal" factor that is set equal to zero. You eventually want to solve for x. So you subtract 1 from both sides and get 2cos(x) = -1. Then to isolate the cos(x) you divide both sides by -2. Now what you have is \[\cos (x) = -\frac{ 1 }{ 2 }\]Are you with me so far?

OpenStudy (anonymous):

im following

OpenStudy (anonymous):

it makes sense when i write it down on paper, its hard when its an online class based on powerpoints

OpenStudy (anonymous):

so basically your breaking it down

OpenStudy (anonymous):

A second version of a Mathematica 9 solution is attached.